A financial planner recommends a plan for a client who has $67,000 to invest.At the end of the presentation, the client asks, "How much will be invested at each rate?" Answer this question using P = 67,000, r1 = 12, r2 = 6.4, and I = 6,662.40.
Answer by ewatrrr(24785)

x the amount at higher interest rate of 12%
.12x + .064(67000-x) = 6662.40
x = (6662.40 - 4288)/.056, amount at 12%
67,000 - x = amount at 6.4%